Lincoln Wastewater Reclamation and Treatment Facility Construction Activity
Work/Activities Performed Last Week:
- Completed forms/rebar for the balance of walkways at Oxidation Ditch 3 – Placed concrete.
- Completed forms/rebar/waterstop at west wall of Oxidation Ditch 3 – Placed concrete.
- Filled new Secondary Clarifier to 2-FT below top-of-wall for pending water leakage test.
- Filled existing Splitter Box addition to 2-FT below top-of-wall for pending water leakage test.
- Backfilled at perimeter of foundation/slab at the Solids Dewatering Building.
- Continued installation of joint sealant at Tertiary Filter and Oxidation Ditch 3.
- Continued sack & patch at interior of Oxidation Ditch 3.
- Continued installation of SLG Gates at existing Splitter Box addition.
Work Scheduled to be Performed this Week:
- Complete prep and start epoxy lining at new Chemical Storage area.
- Start & complete water leakage test of new Secondary Clarifier.
- Start & complete water leakage test of existing Clarifier Splitter Box addition.
- Demo existing wall of Clarifier Splitter Box for new Splitter Box addition water flow.
- Complete prep and start installation of interior epoxy lining at Headworks.
- Re-start structural backfill at perimeter of new RAS Pump Station.
- Pass water leakage test of new Secondary Clarifier and start structural backfill.
- Pass water leakage test of existing Clarifier Splitter Box addition and start structural backfill.
Pipeline Construction Activity
Work/Activities Performed Last Week:
Reach 1 (Sta: 10+00 to 55+00) – Turkey Creek
- This area continues to be used as a staging area for processing soils to be used at other locations on the project. Import soils that are being used on the Wise Road Overland Reach are also being processed at this location.
Reach 1 (Sta: 10+00 to 261+42.90) – Virginiatown Road
- The contractor is moving back to this area. Work will begin on Monday (10/27/14) morning, beginning at station 55+80 and working down-station (westerly).
- The surveyors are staking this area from approximately Sta: 48+00 to 56+00. This will be the portion of the installation that will bring the HDPE pipe from the Auburn Ravine, to the existing end of the pipe at approximately Sta: 55+80.
- Additional work in this area will include a couple of 45 degree welds that have not yet been completed.
Reach 2 (Sta: 10+00 to 109+71.09) – Chili Hill Road
- Trenching operations have been completed up Chili Hill Road to Sta: 91+00. At Sta: 91+00 the HDPE pipe was cut and buried. The excavation was then paved and all equipment demobilized until a future date.
- Two (2) CMPs were cut and set aside for future repair (good condition). These areas have been photographed and documented with the stationing and condition.
- Friday was spent cleaning up the Chili Hill work zone and repairing and replacing SWPPP materials.
- Compaction testing was coordinated with Tim Davis from Blackburn.
- Observation of the end of the day operations; clean-up, paving, placing the plates and trench shoring.
Reach 3 (Sta: 10+00 to 203+71.07) – Wise Road Overland (WRO)
- Work has begun from Sta: 87+00 to 98+00; clearing and grubbing, installation of ESA fencing and access road constructed. Additionally, pipe has been delivered and stockpiled.
- Contractor continues to excavate, lay and backfill pipe on the Shanley and Thompson Overland portion of Reach 3, Sta: 108+70 to Sta: 171+50. This week they installed 635 feet of HDPE pipe, for an average of 127 feet per day.
- 270 feet of trench was excavated, which included the rock area that was blasted between Sta: 139+90 to Sta: 142+65.
- The contractor completed CLSM backfill between Sta. 109+60 to Sta. 108+70. This completes the CLSM backfill down to the 45-degree angle point.
- Between Sta: 108+70 to Sta: 111+35 and between Sta: 118+00 to Sta: 124+00, the contractor has hydro-seeded about an acre of ground.
- The disturbed area between Sta: 152+00 to Sta: 168+00 was fiber mulched preparing for the rain event forecasted for Saturday.
- Between Sta: 88+00 and Sta: 108+15 silt fencing was placed and the overland section was potholed.
Work Schedule to be performed this coming week:
- Processing soils with the power-screen at the Turkey Creek site.
- Continue with trench excavation and pipe installation at Reach 3, S Bar V.
- Continue trench excavation between Sta: 129+25 to Sta: 135+15.
- Continue backfilling and returning to natural condition between Sta: 11+35 to Sta: 122+00 on Reach 3.
- Continue welding and string pipe between Sta: 161+50 and 172+00 in Reach 3.
- Continue to install and monitor the SWPPP best management practices.
SMD-1 Construction Activity
Work/Activities Performed for the identified time period:
- MCC/Electrical Building – DCI – Stripped form panels from slab and started to patch the surfaces where the concrete precast walls will be constructed.
- DCI – Patching leaking areas on the South & West side of Pump Station.
- DCI – Pump Station Wetwell – filling inside of Pump Station with dewatering water, structure was full by Thursday and test started.
- Standby Generator Pad Area – DCI – did minimal cleanup work as the delivery of the generators will now be later in November.
- Auburn Contractors – Excavated and Installed the 4” PVC primary conduit for the PG&E service with sand bedding and had to wait until Friday for the PG&E inspector to review the sand bedding and conduit, then came back later on Friday to inspect the sand shading and the joint trench 2” PVC telephone line that will share the same pole, once installed. The trench split on the East side of MCC / Electrical building with the PG&E conduit rising up into the future transformer location and the telephone conduit rising up to a future pull box required by the phone company.
- DCI – Continued cleaning forms and stacking on pallets with banding for shipping back to EFCO. 90% of the rented forms have been returned to EFCO.
- DCI – Patching Pump Station rock pockets on the South & North side.
- DCI – No Potholing – at SE part of plant to verify the Gas & NID lines that will be crossed during the 28” HDPE force main installation scheduled in approximately two weeks.
- DCI – Additional pipe materials, valves, and pond liner material on palates, were delivered this week along with other specialty items. The yard was reorganized to allow work to be laid out for start of construction at the biofilter site in a couple of weeks.
- SWPPP – DCI added BMP’s for the projected weekend wet weather and started to lay out the silt fence and wattle areas for the 28” forcemain installation on the SMD- 1 project work area.
Work/Activities Planned for this week:
- DCI – The EFCO forms will be removed, cleaned, and palletized from the remaining walls after the water test is complete.
- DCI – The Pump Station exfiltration test was started last Thursday and should be complete this week after repairs to any visible leak areas.
- Camblin steel – No work planned this week.
- DCI – The cement finisher crews will patch the out of tolerance areas for the MCC / Electrical building pad to allow delivery and installation on Wednesday.
- DCI – General site clean-up and moving delivered materials to the NE side of project so DCI can start on Bio-filter work.
- Auburn Contractors – The crews should set the new switchgear equipment in conjunction with the precast walls, scheduled for Wednesday, prior to roof placement.
- No work on Headworks while waiting on DCI pumping plan to be approved or commented on by County and Stantec.
- Possible pothole investigation for possible force main conflicts may continue this week, still looking for the 1” diameter NID water line on SW corner of Plant.